# Antiferromagnetism and Tightly Bound Cooper Pairs Induced by Kinetic Frustration

### v1

@@ C.1 Motivation and methodology
To capture the essential physics of antiferromagnetic polarons as pairing glue, we develop a targeted Hilbert space truncation strategy that selectively incorporates spin fluctuations in the vicinity of doped holes. This strategy enables exact diagonalization of systems large enough to study Cooper pair formation, while preserving physical accuracy for hole kinetics and computational tractability. Our method selectively incorporates quantum fluctuations in regions where they significantly affect hole pairing, while treating distant regions as classical Néel order in Section B.
The physical rationale for this approach stems from the belief of nearsightedness, where fluctuation distant to holes will not affect interactions between holes. Namely, spin fluctuations far from holes contribute nearly identically regardless of hole separation, effectively canceling out in binding energy calculations. Moreover, regions surrounding holes maintaining stronger Néel order due to dominant $t_{2}$ hopping, thus having lower numbers of singlets. By restricting quantum fluctuations to regions where they genuinely impact hole dynamics—and treating more distant areas as a static Néel background—we systematically reduce the Hilbert space size while preserving the essential many-body correlations responsible for pairing. This method eliminates irrelevant background contributions and captures the microscopic mechanisms underlying polaron binding. Similar truncation schemes have proven successful in other strongly correlated contexts, such as electron-phonon coupled systems [56].
@@ C.2 Single-hole basis construction
Our Hilbert space construction begins with a systematic methodology first for the single-hole case, laying the groundwork required to effectively address the complexities inherent in the two-hole problem. balance computational efficiency with physical accuracy through two essential steps.
Firstly, we adopt the classical Néel-ordered state as our initial configuration, characterized by spins uniformly oriented up on one sublattice and down on the other. This choice aligns with our findings that kinetic effects alone inherently stabilize antiferromagnetic correlations, even in the absence of spin exchange interactions. A single hole is then introduced into this ordered background.
Second, we expand the Hilbert space by incorporating controlled kinetic and spin flip processes. We begin by allowing the hole to move via a single nearest-neighbor ($t_{1}$) or next-nearest-neighbor ($t_{2}$) hopping, generating the minimal set of kinetically accessible configurations. After these kinetic moves, we incorporate local spin fluctuations by applying spin-flip operators $S_{i}^{+}S_{j}^{-}+S_{i}^{-}S_{j}^{+}$ with $i,j$ to be the neighboring spin, which capture essential quantum magnetic dynamics around the hole. Here, we restrict spin-flip pairs $(i,j)$ to lie within an $L^{\infty}$ distance of 2 from the hole position. The choice of using $L^{\infty}$ distance is motivated by the coexistence of NN and NNN hopping, and we adopt it consistently throughout our basis truncation scheme. This localization captures the essential physics of hole-induced spin correlations while avoiding exponential growth in configuration space, resulting in a manageable single-hole Hilbert space of 370 states per momentum sector.
Our approach offers a significant improvement over earlier recursive schemes based solely on repeated kinetic propagation [29]. As illustrated in Fig. S4, our method—consisting of one-step hole motion followed by targeted spin fluctuations—successfully captures many high-weight configurations that are inaccessible to recursive approaches. These include states with relatively simple spin textures but nontrivial hole trajectories, which contribute substantially to the true ground state wavefunction yet are overlooked by purely kinetic generation of basis sets.
Figure S4: The configurations with their amplitudes, generated with various basis scheme. Only nonequivalent configurations (with translations, rotations and reflections) with top 30 large amplitude are visualized for each scheme. To compare different scheme the maximum distance between spin flips and holes are set to be 5.
@@ C.3 Two-hole basis construction
Figure S5: (a) Number of fluctuations per kinetically generated configuration. Fixing the range of considered fluctuations tends to reduce the Hilbert space per configuration, leading to effective repulsion, whereas fixing the Hilbert space per relative position (treating holes on the same/different sublattices separately) tends to overestimate the Hilbert space, resulting in effective attraction. (b, c, d) Comparison between ED and DMRG for the hole-hole correlation $P_{i,j}=\langle\bar{n}_{i}\bar{n}_{j}\rangle_{G}$ under parameters $(\lambda,\Delta)=(1,1)$ and $U=10.0$, using the “threshold distance” = 2, enlarging the Hilbert space by fixing the Hilbert space per hole relative positions, and enlarging the Hilbert space by fixing the Hilbert space per kinetic configuration. Fixing the Hilbert space per kinetically generated configuration gives the best match with the DMRG result.
@@
We now extend the Hilbert space construction method developed for the single-hole scenario systematically to construct a two-hole basis, enabling unbiased energy comparisons across diverse hole configurations. Although analogous principles guide the construction, additional treatment is required to control Hilbert space dimensions, avoiding artificial biases during pairing calculations as shown in Fig. S5.
@@
We initiate our two-hole basis construction by placing two holes into a classical Néel-ordered background, ensuring they occupy opposite sublattices to maintain a total $S_{z}=0$. Subsequently, we allow each hole to undergo at most one hopping step, either nearest-neighbor ($t_{1}$) or next-nearest-neighbor ($t_{2}$), to generate fundamental kinetic generated configurations.
@@
Following these kinetic moves, local spin fluctuations are introduced around each hole. Each hole can associate with at most one spin-flip pair, generated by applying the spin-flip operators $S_{i}^{+}S_{j}^{-}+S_{i}^{-}S_{j}^{+}$ on kinetically generated basis states, where sites $(i,j)$ are either nearest or second-nearest neighbors. We apply a maximum distance constraint between each spin-flip pair and its associated hole, ensuring that when holes are sufficiently separated, the resulting state can be decomposed into a product of single-hole states, thereby preventing artificial long-range correlations.
@@
A significant challenge emerges when comparing energies for varying hole separations. Using a fixed maximum distance (hole-spin flip) constraint typically yields a smaller Hilbert space for adjacent holes compared to widely separated ones. Such a difference in subspace dimensions results in artificial biases, as configurations with larger Hilbert spaces naturally achieve lower energies due to increased variational freedom, thus misrepresenting underlying genuine physical interactions.
@@
This straightforward approach, employing a uniform maximum distance (hole-spin flip) constraint, fails to address the overlap of spin-fluctuation configurations when holes are adjacent. Although our construction associates spin flips independently with each hole—effectively creating a product Hilbert space—many configurations perceived as distinct in this product space actually represent identical physical states. For instance, spin-flip pairs around one hole can equivalently emerge from fluctuations near an adjacent hole. Such redundancy significantly reduces the number of accessible states for adjacent holes, artificially elevating their energies and underestimating the true pairing attraction. Physically, this enforced classical ordering in extensive regions near adjacent holes leads to artificially higher energies compared to the true quantum ground state.
@@
Fig. S5 demonstrates this bias quantitatively. We show the average fluctuation-generated Hilbert space per kinetic configuration, defined as the ratio of subspace dimensions with and without adding fluctuations for holes at specific relative positions. When two holes are directly adjacent, the effective Hilbert space size is nearly halved compared to well-separated holes due to configuration overlap (see Fig. S5(a)). Consequently, pairing tendencies become severely underestimated as shown in Fig. S5(b).
To address this limitation, we implement a systematic priority ordering scheme for spin-flip configurations based on hierarchical distance criteria. Initially, we set a threshold distance that ensures independent fluctuations when holes are sufficiently separated. Subsequently, we evaluate distances between hole positions and potential spin-flip sites, prioritizing configurations that minimize the maximum distance, thus favoring compact arrangements. If configurations share the same maximum distance, further ranking criteria based on second and third maximum distances are systematically applied. This hierarchical prioritization consistently favors configurations with smaller overall spatial extents, ensuring unbiased and physically meaningful energy comparisons across hole separations.
We tested two basis enlargement schemes, both including all states with threshold distance less than 3. Relative hole configurations with smaller Hilbert spaces are complemented with new configurations following the priority scheme to reach fixed dimensions. Fig. S5(a) shows that fixing the average Hilbert space dimension per kinetically generated configuration provides the best agreement with DMRG results. Alternative approaches, such as employing simple product spaces or fixing fluctuation-generated Hilbert spaces based solely on hole positions rather than kinetic configurations, yield significantly worse agreement. To account for even-odd effects between holes on same or different sublattices, we fix average Hilbert space sizes separately for these two cases in both enlargement schemes.
@@
We note that our basis construction scheme has limitations at specific region of $(\lambda,\Delta)$ parameter space. Specifically, our approach is optimized for cases where both $\lambda$ and $\Delta$ are simultaneously nonzero, but becomes less reliable when $\lambda=0$ but $\Delta$ is large. In this limit, many kinetically-generated states possess negligible amplitudes due to the absence of $t_{1}$ hopping, which is essential for inter-sublattice hole movement. However, when two holes are positioned nearby, certain kinetically-generated configurations—where both holes move to the opposite sublattice and create a single spin mismatch at their original positions—can still be accessed through spin fluctuations acting on the classical background. Critically, these processes exclusively occur when holes are adjacent, artificially enhancing the pairing attraction in this parameter regime (see Fig. S6(c)).
C.4 Momentum sector selection for Superconducting gap calculation
@@
Figure S7: Momentum sector dependence of hole-hole correlations demonstrates optimal conditions for superconducting gap calculations. We show correlation patterns $P_{i,j}=\langle\bar{n}_{i}\bar{n}_{j}\rangle_{G}$ in different energy eigenstates and momentum sectors, calculated using constrained ED with $(\lambda,\Delta)=(1,1)$ and $U=10t_{1}$. Position $i$ is fixed at the lattice center while $j$ varies across all sites. (a) Ground state at momentum $(0,0)$ showing tightly bound Cooper pairs. (b) First excited state at momentum $(0,0)$ exhibiting hybrid character with both binding and separated-hole components. (c) Ground state at momentum $(0,2\pi/L)$ demonstrating preserved bound pair structure under mixed boundary conditions. (d) First excited state at momentum $(0,2\pi/L)$ showing clean unbound state with minimal pairing correlations. The dramatic difference between excited states in different momentum sectors motivates our choice of $(0,2\pi/L)$ for gap calculations.
@@
We demonstrate that careful selection of the momentum sector plays a critical role in suppressing finite-size effects in superconducting gap calculations. Through a systematic comparison of ground and excited states across different momentum sectors, we identify configurations that best preserve the physical integrity of Cooper pairing while minimizing finite-size artifacts introduced by boundary conditions.
@@
The ground state maintains tightly bound Cooper pairs across different momentum sectors, with characteristic binding extending several lattice spacings. Fig. S7 (a, c) shows that hole-hole correlation patterns remain largely unchanged when varying the momentum sector from $(0,0)$ to $(0,2\pi/L)$. The minimal sensitivity of the ground state correlations to the momentum sector reflects the inherent stability of pair formation and effectively implements mixed boundary conditions—periodic in one direction and antiperiodic in the other—for the relative coordinate of the hole pair.
@@
In contrast, the first excited state shows pronounced momentum-sector dependence. At $k=(0,0)$, we observe a hybrid character displaying both binding correlations and significant probability amplitude for widely separated holes (see Fig. S7 (b)). At $k=(0,2\pi/L)$, however, the excited state transforms into a configuration nearly invisible to the pairing potential, with holes effectively decoupled (see Fig. S7(d)). This latter state achieves slightly lower energy than the hybrid state, allowing a more accurate and unbiased estimate of the superconducting gap.
@@ C.5 Pairing symmetry
We identify the pairing symmetry of the superconducting ground state by systematically analyzing pair–pair correlation functions across different bond orientations. This approach allows us to distinguish between different pairing—such as $s$-wave and $d_{x^{2}-y^{2}}$-wave symmetries—based on how pair correlations transform under lattice rotations and reflections.
@@
$C_{i,j}^{s,\|}$ $=$ $\left\langle\Psi_{G}\left|\hat{\Delta}_{i,i+\hat{\mathbf{e}}_{y}}^{s}\left(\hat{\Delta}_{j,j+\hat{\mathbf{e}}_{y}}^{s}\right)^{\dagger}\right|\Psi_{G}\right\rangle,$ (19)
$C_{i,j}^{s,\perp}$ $=$ $\left\langle\Psi_{G}\left|\hat{\Delta}_{i,i+\hat{\mathbf{e}}_{y}}^{s}\left(\hat{\Delta}_{j,j+\hat{\mathbf{e}}_{x}}^{s}\right)^{\dagger}\right|\Psi_{G}\right\rangle$ (20)
Here we fix the reference pair $i,i+\hat{\mathbf{e}}_{y}$ at the lattice center and vary over all horizontal and vertical bonds that do not overlap with the central pair creation position.
@@
Figure S8: Pair-pair correlations confirm $d_{x^{2}-y^{2}}$ pairing symmetry through characteristic sign changes under rotation. We calculate $\left\langle\Psi_{G}\left|\hat{\Delta}_{i,i+\hat{\mathbf{e}}_{y}}^{s}\left(\hat{\Delta}_{j,k}^{s}\right)^{\dagger}\right|\Psi_{G}\right\rangle$ with reference pair fixed at lattice center ($i$) and test pairs at all nearest-neighbor bonds $(j,k)$. The sign change between horizontal pairs ($k=j+\hat{\mathbf{e}}_{x}$) and vertical pairs ($k=j+\hat{\mathbf{e}}_{y}$) confirms $d_{x^{2}-y^{2}}$ symmetry. (a) DMRG calculation on 8-leg cylinder with $(\lambda,\Delta)=(1,1)$ and $U=10t_{1}$. (b) Constrained ED calculation on $16\times 16$ periodic cluster with identical parameters, scaled by factor of 4 to account for different pair densities.

### v2

@@ C.1 Motivation and methodology
To capture the essential physics of antiferromagnetic polarons as pairing glue, we develop a targeted Hilbert space truncation strategy that selectively incorporates spin fluctuations in the vicinity of doped holes. This strategy enables exact diagonalization of systems large enough to study Cooper pair formation, while preserving physical accuracy for hole kinetics and computational tractability. In the spirit of the original method introduced by Trugman [34, 51], our method selectively incorporates quantum fluctuations in regions where they significantly affect hole pairing, while treating distant regions as classical Néel order in Section B.
The physical rationale behind this approach is rooted in the principle of nearsightedness, which holds that distant background spin fluctuations have a negligible effect on the interactions between holes. Specifically, spin fluctuations far from the holes contribute nearly identically regardless of the hole separation, and therefore cancel out in binding energy calculations. Moreover, the spin background retains strong Néel order due to the dominant $t_{2}$ hopping and antiferromagnetic exchange interactions.
By restricting quantum fluctuations to regions where they meaningfully influence hole dynamics—and treating more distant areas as a static Néel background—we systematically reduce the Hilbert space size while preserving the essential many-body correlations responsible for pairing. This approach eliminates irrelevant background contributions and captures the microscopic mechanisms underlying polaron binding. Similar truncation schemes have proven effective in other strongly correlated systems, such as electron-phonon coupled models [60].
@@ C.2 Single-hole basis construction
Our Hilbert space construction begins with a systematic methodology for the single-hole case, which lays the foundation for addressing the complexities of the two-hole problem. The goal is to balance computational efficiency with physical accuracy through two essential steps.
First, we adopt the classical Néel-ordered state as the initial configuration, with spins aligned up on one sublattice and down on the other. This choice is motivated by our finding that kinetic effects alone can stabilize antiferromagnetic correlations, even in the absence of explicit spin exchange interactions. A single hole is then introduced into this ordered background.
Second, we expand the Hilbert space by incorporating controlled kinetic and spin-flip processes. We begin by allowing the hole to move through a single nearest-neighbor ($t_{1}$) or next-nearest-neighbor ($t_{2}$) hopping, thereby generating the minimal set of kinetically accessible configurations. These kinetic processes are recursively applied to all charge/spin configurations up to ${D_{k}}$ times. In our current approach, we set ${D_{k}}=1$, whereas previous studies typically use ${D_{k}}\approx 3\to 5$. After these kinetic moves, we incorporate local spin fluctuations by applying spin-flip operators of the form $S_{i}^{+}S_{j}^{-}+S_{i}^{-}S_{j}^{+}$, where $(i,j)$ are neighboring spins. These operators capture essential quantum magnetic dynamics in the vicinity of the hole. We restrict the spin-flip pairs to lie within the “threshold distance”, an $L^{\infty}$ distance ${D_{f}}$ from the hole position. The use of the $L^{\infty}$ metric is motivated by the coexistence of NN and NNN hopping, and it is consistently adopted throughout our basis truncation scheme. This localization strategy captures the key physics of hole-induced spin correlations while avoiding exponential growth in the Hilbert space, resulting in a manageable single-hole basis of 370 states per momentum sector with ${D_{f}}=2$.
Our approach represents a significant improvement over earlier recursive schemes based solely on repeated kinetic propagation [34]. As illustrated in Fig. S4, our method—which combines one-step hole motion with targeted spin fluctuations—successfully captures many high-weight configurations that are inaccessible to recursive approaches. These include states with relatively simple spin textures but nontrivial hole trajectories, which contribute substantially to the ground state wavefunction and are missed when the basis is generated purely through kinetic processes.
Figure S4: Configurations with their amplitudes generated with various basis schemes. Only nonequivalent configurations (up to translations, rotations and reflections) with top 30 large amplitude are visualized for each scheme. The maximum distance between spin flips and holes for the improved method is set to be ${D_{f}}=5$. We compare it against repeated kinetic propagation method with ${D_{f}}=0$ , where the value ${D_{k}}$ indicates the number of kinetic propagation recursively applied in vertical, horizontal and diagonal directions to the classical Néel spin configurations.
@@ C.3 Two-hole basis construction
Figure S5: (a) Number of magnetic states per hole motion generated configuration of two holes. On one hand, fixing the range of magnetic fluctuations per individual hole tends to reduce the Hilbert when the distance between holes is smaller than the allowed range of fluctuations ${D_{f}}$ (because the same state is generated in more than one way), leading to effective repulsion. On the other hand, fixing the Hilbert space dimension for arbitrary relative position (treating holes on the same/different sublattices separately) biases the system by generating an artificial attractive potential. (b, c, d) Comparison between ED and DMRG for the hole-hole correlation function $P_{i,j}=\langle\bar{n}_{i}\bar{n}_{j}\rangle_{G}$ for $(\lambda,\Delta)=(1,1)$ and $U=10.0$, using i) the “threshold distance” ${D_{f}}$= 2, enlarging the Hilbert space dimension ii) by fixing the Hilbert space dimension for any relative position of the two holes, and iii) by making the Hilbert space dimension for each relative position between the two holes proportional to the number of magnetic states generated by hole motion (see main text). The third option gives the best match with the DMRG result.
@@
We now systematically extend the Hilbert space construction method developed for the single-hole case to the two-hole scenario, enabling unbiased energy comparisons across a wide range of hole configurations. While the construction follows similar guiding principles, additional care is required to control the Hilbert space dimension and avoid artificial biases in pairing energy estimates, as illustrated in Fig. S5.
@@
We begin by placing two holes into a classical Néel-ordered background, ensuring that they occupy opposite sublattices to maintain total $S_{z}=0$. Each hole is then allowed to undergo a single hopping step, either nearest-neighbor ($t_{1}$) or next-nearest-neighbor ($t_{2}$), thereby generating the minimal set of kinetically accessible configurations. With the language of single hole case, this corresponds to ${D_{k}}=1$ for both holes.
@@
Next, we introduce local spin fluctuations around each hole. Each hole may be associated with at most one spin-flip pair, created by applying the bond operator $S_{i}^{+}S_{j}^{-}+S_{i}^{-}S_{j}^{+}$ to the kinetically generated basis states, where $(i,j)$ are either nearest or second-nearest neighbors. We impose a maximum $L^{\infty}$ distance ${D_{f}}$ between the spin-flip pair and its associated hole to ensure that the resulting configurations can be decomposed into a product of two independent single-hole states when the holes are far apart. This constraint eliminates unphysical long-range correlations and ensures that the spin fluctuations are localized where they meaningfully impact hole dynamics.
@@
A key challenge arises when comparing energies for different hole separations. Using a fixed maximum spin-flip distance typically yields a smaller Hilbert space dimension for adjacent holes compared to well-separated ones. This disparity introduces a systematic bias: configurations with larger Hilbert space of magnetic fluctuations benefit from greater variational freedom and thus artificially lower energies, obscuring the true physical nature of hole pairing.
@@
This issue is compounded by the fact that our construction treats spin flips as independently associated with each hole, effectively forming a product basis. When the holes are close together, many configurations that appear distinct in this product space actually represent the same physical state due to overlap in the fluctuation regions. For example, spin flips assigned to one hole may equivalently be attributed to its neighbor. This redundancy reduces the effective Hilbert space for adjacent-hole configurations, artificially raising their energy and underestimating the pairing interaction. Physically, this manifests as an enforced classical ordering near the holes, which is energetically unfavorable relative to the true quantum ground state.
@@
Fig. S5 quantifies this bias. In panel (a), we show the average fluctuation-generated Hilbert space per kinetic configuration, defined as the ratio of the fixed hole position subspace dimensions with and without spin fluctuations for fixed hole separations. When two holes are adjacent, the effective Hilbert space is nearly halved compared to the well-separated case due to configuration overlap. As a result, pairing tendencies are severely underestimated, as shown in panel (b).
To mitigate this bias, we implement a hierarchical prioritization scheme for generating spin-flip configurations. We begin by including all magnetic configurations generated with ${D_{k}}=1$ and ${D_{f}}=2$. As previously explained, when the two holes located at positions $\bm{r}_{i}$ and $\bm{r}_{j}$ are close, the ratio between the number of total magnetic configurations, $N_{\text{fl}}(\rho)=N(D_{k}=1,D_{f}=2,\rho)$ with $\rho=|\bm{r}_{i}-\bm{r}_{j}|$ and the number of kinetically generated configurations, $N_{\text{kin}}(\rho)=N(D_{k}=1,D_{f}=0,\rho)$, becomes significantly smaller. This motivates the increase of configurations due to magnetic fluctuations for small $\rho$ by increasing $D_{f}$.
The extension is guided by a prioritization scheme that favors configurations in which the spin-flip pair is as close as possible to its associated hole. When multiple configurations share the same maximum hole spin-flip distance, we apply secondary and tertiary criteria based on the next-largest distances. All configurations meeting these criteria are included, and any configurations tied under the same ranking are retained to preserve symmetry. We continue this process until adding further configurations would increase the maximal ratio $\max_{i,j}N_{\text{fl}}(\rho)/N_{\text{kin}}(\rho)$. This maximum ratio is fixed based on configurations with $D_{k}=1,D_{f}=2$, and we obtain the maximum separately for hole pairs on the same and different sublattices, acknowledging that sublattice effects remains relevant when the holes are far apart. This hierarchical approach favors compact configurations to address Hilbert space discrepancies, ensuring that the strongest hole-magnon interactions are fully captured.
We tested two basis enlargement schemes, one following the method illustrated above, and the other do in a much naive way, fixing $\max_{i,j}N_{\text{fl}}(\rho)$ with the same prioritization scheme. Fig. S5 shows that our final scheme yields the best agreement with DMRG results. In contrast, simpler approaches, such as fixing ${D_{k}}=1$ and ${D_{f}}=2$ or fixing the Hilbert space dimension $N_{\text{fl}}(\rho)$ for any relative position of the two holes, lead to significantly poorer agreement.
@@
We note that the accuracy of our basis construction scheme is limited in certain regions of the $(\lambda,\Delta)$ parameter space. In particular, the approach is optimized for regimes where both $\lambda$ and $\Delta$ are simultaneously nonzero. When $\lambda=0$ but $\Delta$ is large, the reliability decreases. In this limit, many kinetically generated states carry negligible amplitude due to the absence of $t_{1}$ hopping, which is essential for inter-sublattice hole motion. Nevertheless, for nearby hole configurations, it is still possible to generate kinetically generated states via spin fluctuations acting on the classical background—specifically, which gives identical configurations as both holes move to the opposite sublattice and create a single spin mismatch at their original positions. These processes occur exclusively when the holes are adjacent and can artificially enhance pairing in this regime, as shown in Fig. S6(c).
C.4 Momentum sector for superconducting gap
@@
Figure S7: Momentum sector dependence of hole-hole correlations demonstrates optimal conditions for superconducting gap calculations. We show correlation patterns $P_{i,j}=\langle\Psi_{G}|\bar{n}_{i}\bar{n}_{j}|\Psi_{G}\rangle$ in different energy eigenstates and momentum sectors, calculated using constrained ED with $(\lambda,\Delta)=(1,1)$ and $U=10t_{1}$. Position $i$ is fixed at the lattice center while $j$ varies across all sites. (a) Ground state at momentum $(0,0)$ showing tightly bound Cooper pairs. (b) First excited state at momentum $(0,0)$ exhibiting hybrid character with both binding and separated-hole components. (c) Ground state at momentum $(0,2\pi/L)$ demonstrating preserved bound pair structure under mixed boundary conditions. (d) First excited state at momentum $(0,2\pi/L)$ showing clean unbound state with minimal pairing correlations. The dramatic difference between excited states in different momentum sectors motivates our choice of $(0,2\pi/L)$ for gap calculations.
@@
We demonstrate that the careful selection of momentum sectors plays a crucial role in suppressing finite-size effects in superconducting gap calculations. By systematically comparing ground and excited states across different momentum sectors, we identify configurations that best preserve the physical nature of Cooper pairing while minimizing artifacts induced by boundary conditions.
@@
The ground state exhibits tightly bound Cooper pairs across all momentum sectors, with characteristic coherence lengths extending over several lattice spacings. As shown in Fig. S7(a, c), the hole-hole correlation patterns remain largely unchanged when the momentum is varied from $(0,0)$ to $(0,2\pi/L)$. This insensitivity highlights the robustness of pair formation and effectively mimics mixed boundary conditions—periodic in one direction and antiperiodic in the other—for the relative coordinate of the hole pair.
@@
In contrast, the first excited state shows a pronounced sensitivity to the choice of momentum sector. At $\bm{k}=(0,0)$, it exhibits a hybrid character, combining signatures of bound-pair correlations with a significant probability amplitude for widely separated holes [Fig. S7(b)]. In the $\bm{k}=(0,2\pi/L)$ sector, however, the excited state evolves into a nearly unbound configuration, effectively decoupled from the pairing potential [Fig. S7(d)]. This unbound state attains slightly lower energy than the hybrid configuration, providing a more accurate and less biased estimate of the superconducting gap.
@@ C.5 Pairing symmetry
We identify the pairing symmetry of the superconducting ground state by systematically analyzing pair–pair correlation functions across different bond orientations. This approach allows us to distinguish between different pairing symmetries —such as $s$-wave and $d_{x^{2}-y^{2}}$-wave— based on how pair correlations transform under lattice rotations and reflections.
@@
$C_{i,j}^{s,\|}$ $=$ $\left\langle\Psi_{G}\left|\hat{\Delta}_{i,i+\hat{\bm{e}}_{y}}^{s}\left(\hat{\Delta}_{j,j+\hat{\bm{e}}_{y}}^{s}\right)^{\dagger}\right|\Psi_{G}\right\rangle,$ (20)
$C_{i,j}^{s,\perp}$ $=$ $\left\langle\Psi_{G}\left|\hat{\Delta}_{i,i+\hat{\bm{e}}_{y}}^{s}\left(\hat{\Delta}_{j,j+\hat{\bm{e}}_{x}}^{s}\right)^{\dagger}\right|\Psi_{G}\right\rangle$ (21)
Here we fix the reference pair $i,i+\hat{\bm{e}}_{y}$ at the lattice center and vary over all horizontal and vertical bonds that do not overlap with the central pair creation position.
@@
Figure S8: Pair-pair correlations confirm $d_{x^{2}-y^{2}}$ pairing symmetry through characteristic sign changes under rotation. We calculate $\left\langle\Psi_{G}\left|\hat{\Delta}_{i,i+\hat{\bm{e}}_{y}}^{s}\left(\hat{\Delta}_{j,k}^{s}\right)^{\dagger}\right|\Psi_{G}\right\rangle$ with reference pair fixed at lattice center ($i$) and test pairs at all nearest-neighbor bonds $(j,k)$. The negative relative sign between horizontal pairs ($k=j+\hat{\bm{e}}_{x}$) and vertical pairs ($k=j+\hat{\bm{e}}_{y}$) confirms $d_{x^{2}-y^{2}}$ symmetry. (a) DMRG calculation on 8-leg cylinder with $(\lambda,\Delta)=(1,1)$ and $U=10t_{1}$. (b) Constrained ED calculation on $16\times 16$ periodic cluster with identical parameters, scaled by factor of 4 to account for different pair densities.